OpenStudy (anonymous):

give example of eq, one real variable x such that it is not linear and it has (a)no solution (b)unique solution (C)two solution only (d) infinite solution

OpenStudy (anonymous):

(a)no solution \(e^x=-1\) ..... \(1+x^2=0\) ......... \(\sqrt{x}+\sqrt{x+16}=3\) (b)unique solution \(x^2=0\) ............. \(\sqrt{x}=5\) (c)two solution only \((x-1)(x-2)=0\) .......... (d) infinite solution \(\sin x=0\)

OpenStudy (anonymous):

please give grahical approch to this problem

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

for example \(e^x=-1\)  as u can see always \(e^x >0\) so its impossible to find a \(x\) for which \(e^x=-1\)